# Multiple Imputation for Nonresponse in Complex Surveys Using Design Weights and Auxiliary Margins

### v3

@@ Appendix A Simulation Study Parameters
This section presents the parameter values used in the simulation studies of Section 4, which are identical to those in Yang and Reiter (2025). We set $\nu_{u}=-1.2$; $(\omega_{10},\omega_{11})=(0.06,-0.0002)$ and $\theta_{1}\in(-2,-0.5)$; $(\omega_{20},\omega_{21})=(0.2,0.4)$ and $\theta_{2}=-2$; $(\omega_{30},\omega_{31},\omega_{32})=(0.2,0.3,0.1)$; $(\omega_{40},\dots,\omega_{43})=(0.2,0.4,0.4,0.1)$; $(\omega_{50},\dots,\omega_{54})=(0.4,1.2,-0.9,0.1,0.2)$ and $\sigma_{5}=0.5$; $(\omega_{60},\dots,\omega_{65})=(0.4,1.2,-0.9,0.1,-0.1,0.1)$ and $\sigma_{6}=0.5$; $(\phi_{j0},\dots,\phi_{j5})=(-1.4,0.1,0.1,0.1,0.1,0)$ for $j\in\{2,3,4\}$,; and, $(\phi_{60},\dots,\phi_{65})$$=(-1.4,0.1,0.1,0.1,0.1,0.1)$.

### v4

@@ Appendix A Simulation Parameters
The parameter values used in the simulations of Section 4 are identical to those in Yang and Reiter (2025). We set $\nu_{u}=-1.2$; $(\omega_{10},\omega_{11})=(0.06,-0.0002)$ and $\theta_{1}\in(-2,-0.5)$; $(\omega_{20},\omega_{21})=(0.2,0.4)$ and $\theta_{2}=-2$; $(\omega_{30},\omega_{31},\omega_{32})=(0.2,0.3,0.1)$; $(\omega_{40},\dots,\omega_{43})=(0.2,0.4,0.4,0.1)$; $(\omega_{50},\dots,\omega_{54})=(0.4,1.2,-0.9,0.1,0.2)$ and $\sigma_{5}=0.5$; $(\omega_{60},\dots,\omega_{65})=(0.4,1.2,-0.9,0.1,-0.1,0.1)$ and $\sigma_{6}=0.5$; $(\phi_{j0},\dots,\phi_{j5})=(-1.4,0.1,0.1,0.1,0.1,0)$ for $j\in\{2,3,4\}$,; and, $(\phi_{60},\dots,\phi_{65})$$=(-1.4,0.1,0.1,0.1,0.1,0.1)$.